# DiReCT: Diagnostic Reasoning for Clinical Notes via Large Language Models

### v2

@@ 5.2 Results
Table 5: Consistency of automated evaluation metrics with human judgments.
Model Observation Rationalization
LLama3 8B 0.887 0.801
GPT-4 turbo 0.902 0.836
@@
Reliability of automatic evaluation. We randomly pick out 100 samples from DiReCT and their prediction by GPT-4 over the task with $\mathcal{G}$ to assess the consistency of our automated metrics to evaluate the observational and explanatory performance in Section 3.3 to human judgments. Three physicians joined this experiment. For each prediction $\hat{o}\in\hat{\mathcal{O}}$, they are asked to find a similar observation in ground truth $\mathcal{O}$. For explanatory metrics, they verify if each prediction $\hat{z}\in\hat{\mathcal{E}}$ for $\hat{o}\in\hat{\mathcal{O}}$ align with ground-truth $z\in\mathcal{E}$ corresponding to $o$. A prediction and a ground truth are deemed aligned for both assessments if at least two specialists agree. We compare LLama3’s and GPT-4’s judgments to explore if there is a gap between these LLMs. As summarized in Table 5, GPT-4 achieves the best results, with LLama3 8B also displaying a similar performance. From these results, we argue that our automated evaluation metrics are consistent with human judgments, and LLama3 is sufficient for this evaluation, allowing the cost-efficient option.
@@
A prediction example. Figure 6 shows a sample generated by GPT-4. The ground-truth PDD of the input clinical note is Hemorrhagic Stroke. In this figure, purple, orange, and red indicate explanations only in the ground truth, only in prediction, and common in both, respectively; therefore, red is a successful prediction of an explanation, while purple and orange are a false negative and false positive. GPT-4 treats the observation of aurosis fugax as the criteria for diagnosing Ischemic Stroke. However, this observation only supports Suspected Stroke. Conversely, observation thalamic hematoma, which is the key indicator of Hemorrhagic Stroke, is regarded as a less important clue. Such observation-diagnosis correspondence errors lead to the model’s misdiagnosis. More samples are available in the supplementary material.
@@ 6 Conclusion and Limitations
Limitations. DiReCT encompasses only a subset of disease categories and considers only one PDD, omitting the inter-diagnostic relationships due to their complexity—a significant challenge even for human doctors. Additionally, our baseline may not use optimal prompts, chain-of-thought reasoning, or address issues related to hallucinations in task responses. Our dataset is solely intended for model evaluation but not for use in clinical environments. The use of the diagnostic knowledge graph is also limited to serving merely as a part of the input. Future work will focus on constructing a more comprehensive disease dataset and developing an extensive diagnostic knowledge graph.
@@
This work was supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. This work is also supported by JSPS KAKENHI 24K20795 and Dalian Haichuang Project for Advanced Talents.
@@ A.1 Data Statistics
Cardiology Acute Coronary Syndromes 65 6 3 [Byrne et al., 2024]
@@
Gastroenterology Gastritis 27 5 3 [Shah et al., 2021]
@@
Peptic Ulcer Disease 28 3 2 [Kavitt et al., 2019]
@@
Migraine 4 3 2 [Lipton et al., 2001]
Multiple Sclerosis 27 6 4 [Lublin, 2005]
@@
Pulmonology Asthma 13 7 5 [Qaseem et al., 2011]
@@
Pneumonia 20 4 2 [Olson and Davis, 2020]
@@
Endocrinology Adrenal Insufficiency 20 4 3 [Charmandari et al., 2014]
@@
Pituitary 12 4 3 [Tritos and Miller, 2023]
@@
A.2 Structure of Knowledge Graph
@@
The entire knowledge graph, denoted as $\mathcal{K}$, is stored in separate JSON files, each corresponding to a specific disease category $i$ as $\mathcal{K}_{i}$. Each $\mathcal{K}_{i}$ comprises a procedural graph $\mathcal{G}_{i}$ and the corresponding premise $p$ for each disease. As illustrated in Figure 7, the procedural graph $\mathcal{G}_{i}$ is stored under the key "Diagnostic" in a dictionary structure. A key with an empty list as its value indicates a leaf diagnostic node as $d^{\star}$. The premise for each disease is saved under the key of "Knowledge" with the corresponding disease name as an index. For all the root nodes (e.g., Suspected Heart Failure), we further divide the premise into "Risk Factors", "Symptoms", and "Signs". Note that each premise is separated by ";".
@@ B.1 Prompt Settings
In this section, we demonstrate the prompt we used for each module (From Table 7-9 for narrowing-down, perception, and reasoning module, respectively).
@@
In Table 9, {disease} is the disease category name and {disease_option} is consisted by the children nodes $\{d_{n}\}_{n}$. Similarly, the premise on the blue is only available for the $\mathcal{k}$ setting. It provides the premise that are criteria for the diagnosis of each children node. {observation} is the extracted $\hat{\mathcal{O}}$ in previous step. We provide all the prompts and the complete implementation code on GitHub.
B.2 Details of Automatic Evaluation
@@
B.3 Prediction Samples
Figure 9 and 10 shows two sample generated by GPT-4. The ground-truth PDD of the input clinical note is Gastroesophageal Reflux Disease (GERD) and Heart Failure (HF). In these figure, purple, orange, and red indicate explanations only in the ground truth, only in prediction, and common in both, respectively; therefore, red is a successful prediction of an explanation, while purple and orange are a false negative and false positive.
@@ B.3 Prediction Samples
In Figure 9, we can observe that GPT-4 can find the key observation for the diagnosis of GERD, which is consistent with human in both observation and rationale. However, it still lacks the ability to identify all observations and establish accurate relationships for diseases. In Figure 10, the model’s predictions do not align well with those of a human doctor. Key observations, such as the relationships between BNP and LVEF, are incorrectly identified, leading to a final misdiagnosis.
B.4 Experiments for No Extra Knowledge
@@
We demonstrate the prompt used for $\mathcal{D}^{\star}$ and no knowledge settings in Table 12 and Table 13, respectively. {note} is the text of whole clinical note and {note} in Table 12 is the name of all leaf node $\mathcal{D}^{\star}$.
B.5 Experimental Settings

### v3

@@ 5.2 Results
Diagnostic reasoning under conditions of incomplete observation. In real-world scenarios, doctors often have to make diagnoses based on incomplete information. To explore this, we conducted experiments on the 73 amended cases which originally lack observation to the final diagnosis (refer to supplementary for detailed introduction of amended data point). One set of experiments used the unmodified original notes, labeled as "Original," while the other set used notes with added observations labeled as "Amended." We tested three models—Llama3 70B, GPT-3.5-turbo, and GPT-4 turbo—under two settings: one with only the procedural graph $\mathcal{G}$ and the other with the complete knowledge graph $\mathcal{K}$. The results are presented in Tables 5 and 6. We can observe that in both $\mathcal{G}$ and $\mathcal{K}$ settings, the performance on the Amended data was consistently better across all metrics compared to the Original data. This suggests that even a single added observation can significantly impact the model’s diagnostic reasoning.
Table 5: Amendment ablation study using $\mathcal{G}$.
Diagnosis Observation Explanation
@@
Original LLama3 70B 0.547 0.273 0.225±${}_{\pm\text{0.143}}$ 0.472±${}_{\pm\text{0.144}}$ 0.253±${}_{\pm\text{0.138}}$ 0.216±${}_{\pm\text{0.271}}$ 0.073±${}_{\pm\text{0.087}}$
GPT-3.5 turbo 0.507 0.273 0.393±${}_{\pm\text{0.216}}$ 0.355±${}_{\pm\text{0.174}}$ 0.278±${}_{\pm\text{0.151}}$ 0.207±${}_{\pm\text{0.305}}$ 0.062±${}_{\pm\text{0.093}}$
GPT-4 turbo 0.616 0.328 0.446±${}_{\pm\text{0.211}}$ 0.418±${}_{\pm\text{0.164}}$ 0.340±${}_{\pm\text{0.178}}$ 0.242±${}_{\pm\text{0.324}}$ 0.098±${}_{\pm\text{0.137}}$
Amended LLama3 70B 0.698 0.534 0.250±${}_{\pm\text{0.173}}$ 0.507±${}_{\pm\text{0.134}}$ 0.240±${}_{\pm\text{0.129}}$ 0.296±${}_{\pm\text{0.354}}$ 0.133±${}_{\pm\text{0.142}}$
GPT-3.5 turbo 0.671 0.411 0.487±${}_{\pm\text{0.206}}$ 0.351±${}_{\pm\text{0.152}}$ 0.310±${}_{\pm\text{0.145}}$ 0.272±${}_{\pm\text{0.321}}$ 0.092±${}_{\pm\text{0.118}}$
GPT-4 turbo 0.726 0.547 0.546±${}_{\pm\text{0.184}}$ 0.465±${}_{\pm\text{0.148}}$ 0.412±${}_{\pm\text{0.171}}$ 0.391±${}_{\pm\text{0.374}}$ 0.180±${}_{\pm\text{0.186}}$
Table 6: Amendment ablation study using $\mathcal{K}$.
Diagnosis Observation Explanation
[…]
Original LLama3 70B 0.575 0.219 0.109±${}_{\pm\text{0.233}}$ 0.443±${}_{\pm\text{0.171}}$ 0.203±${}_{\pm\text{0.186}}$ 0.304±${}_{\pm\text{0.388}}$ 0.114±${}_{\pm\text{0.135}}$
GPT-3.5 turbo 0.548 0.233 0.293±${}_{\pm\text{0.243}}$ 0.218±${}_{\pm\text{0.198}}$ 0.184±${}_{\pm\text{0.166}}$ 0.251±${}_{\pm\text{0.357}}$ 0.072±${}_{\pm\text{0.106}}$
GPT-4 turbo 0.616 0.260 0.452±${}_{\pm\text{0.241}}$ 0.410±${}_{\pm\text{0.211}}$ 0.349±${}_{\pm\text{0.223}}$ 0.467±${}_{\pm\text{0.437}}$ 0.220±${}_{\pm\text{0.256}}$
Amended LLama3 70B 0.685 0.537 0.261±${}_{\pm\text{0.195}}$ 0.493±${}_{\pm\text{0.230}}$ 0.277±${}_{\pm\text{0.171}}$ 0.452±${}_{\pm\text{0.407}}$ 0.185±${}_{\pm\text{0.194}}$
GPT-3.5 turbo 0.657 0.465 0.390±${}_{\pm\text{0.227}}$ 0.272±${}_{\pm\text{0.194}}$ 0.232±${}_{\pm\text{0.156}}$ 0.401±${}_{\pm\text{0.394}}$ 0.127±${}_{\pm\text{0.145}}$
GPT-4 turbo 0.712 0.589 0.534±${}_{\pm\text{0.214}}$ 0.452±${}_{\pm\text{0.180}}$ 0.401±${}_{\pm\text{0.201}}$ 0.607±${}_{\pm\text{0.442}}$ 0.286±${}_{\pm\text{0.258}}$
[…]
Reliability of automatic evaluation. We randomly pick out 100 samples from DiReCT and their prediction by GPT-4 over the task with $\mathcal{G}$ to assess the consistency of our automated metrics to evaluate the observational and explanatory performance in Section 3.3 to human judgments.
@@
Table 7: Consistency of automated evaluation with human judgments. Evaluated by mean and confidence interval (CI).
Observation Rationalization
Model Mean 95% CI Mean 95% CI
LLama3 8B 0.887 $0.844\sim 0.878$ 0.835 $0.759\sim 0.818$
GPT-4 turbo 0.902 $0.830\sim 0.863$ 0.876 $0.798\sim 0.853$
@@
Three physicians joined this experiment. For each prediction $\hat{o}\in\hat{\mathcal{O}}$, they are asked to find a similar observation in ground truth $\mathcal{O}$. For explanatory metrics, they verify if each prediction $\hat{z}\in\hat{\mathcal{E}}$ for $\hat{o}\in\hat{\mathcal{O}}$ align with ground-truth $z\in\mathcal{E}$ corresponding to $o$. A prediction and a ground truth are deemed aligned for both assessments if at least two specialists agree. We compare LLama3’s and GPT-4’s judgments to explore if there is a gap between these LLMs. As summarized in Table 7, GPT-4 achieves the best results, with LLama3 8B also displaying a similar performance. From these results, we argue that our automated evaluation metrics are consistent with human judgments, and LLama3 is sufficient for this evaluation, allowing the cost-efficient option. Detailed analysis is available in the supplementary material.
@@
Prediction examples. Figure 6 shows a sample generated by GPT-4. The ground-truth PDD of the input clinical note is Hemorrhagic Stroke. In this figure, purple, orange, and red indicate explanations only in the ground truth, only in prediction, and common in both, respectively; therefore, red is a successful prediction of an explanation, while purple and orange are a false negative and false positive. GPT-4 treats the observation of aurosis fugax as the criteria for diagnosing Ischemic Stroke. However, this observation only supports Suspected Stroke. Conversely, observation thalamic hematoma, which is the key indicator of Hemorrhagic Stroke, is regarded as a less important clue. Such observation-diagnosis correspondence errors lead to the model’s misdiagnosis. In Figure 7, we can observe that GPT-4 can find the key observation for the diagnosis of GERD, which is consistent with human in both observation and rationale. However, it still lacks the ability to identify all observations. More samples are available in the supplementary material.
@@ 6 Conclusion and Limitations
Limitations. DiReCT encompasses only a subset of disease categories and considers only one PDD, omitting the inter-diagnostic relationships due to their complexity—a significant challenge even for human doctors. Additionally, our baseline may not use optimal prompts or address issues related to hallucinations in task responses. Our dataset is solely intended for model evaluation but not for use in clinical environments. The use of the diagnostic knowledge graph is also limited to serving merely as a part of the input and once a knowledge graph is provided, the focus shifts to whether the LLM follows the graph’s rules well (refer to supplementary). Future work will focus on constructing a more comprehensive disease dataset and developing an extensive diagnostic knowledge graph.
@@
This work was supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. This work is also supported by JST ACT-X Grant Number JPMJAX24C8, JSPS KAKENHI No. 24K20795 and No. JP23H00497, CREST Grant No. JPMJCR20D3, JST FOREST Grant No. JPMJFR216O, and Dalian Haichuang Project for Advanced Talents.
@@ A.1 Data Statistics
Cardiology Acute Coronary Syndromes 65 6 3 [Byrne et al., 2024, Kitaoka et al., 2020]
@@
Gastroenterology Gastritis 27 5 3 [Shah et al., 2021, of Gastroenterology et al., 2023, Banks et al., 2019, Chow et al., 2010]
@@
Peptic Ulcer Disease 28 3 2 [Kavitt et al., 2019, Tarasconi et al., 2020]
@@
Migraine 4 3 2 [Lipton et al., 2001, Eigenbrodt et al., 2021, Headache Classification Committee of the International Headache Society (2018), IHS]
Multiple Sclerosis 27 6 4 [Lublin, 2005, Brownlee et al., 2017]
@@
Pulmonology Asthma 13 7 5 [Qaseem et al., 2011, Bateman et al., 2007, Baos et al., 2018]
@@
Pneumonia 20 4 2 [Olson and Davis, 2020, RECOMMENDATIONS, 2012, Niederman et al., 2001]
@@
Endocrinology Adrenal Insufficiency 20 4 3 [Charmandari et al., 2014, Yanase et al., 2016, Bornstein et al., 2016]
@@
Pituitary 12 4 3 [Tritos and Miller, 2023, Drummond et al., 2019, Cooper and Melmed, 2012, Mayson and Snyder, 2014]
@@
Notation Description Notation Description
$R$ The whole content of input note. $U$ The narrowing-down module.
$r$ One data section of the input note. $W$ The perception module.
$\mathcal{D}$ Disease collection. $V$ The reasoning module.
$\mathcal{D}^{\star}$ PDD collection. $\{d_{n}\}$ Collection of children diagnosis.
$\mathcal{D}^{\star}_{i}$ PDD collection for disease $i$. $Acc^{diag}$ Diagnosis accuracy for $d^{\star}$.
$d^{\star}$ A PDD disease. $Acc^{cat}$ Diagnosis accuracy for category.
$d_{t}$ Diagnosis at $t$-th iteration. $Obs^{pre}$ Precision of observation.
$d$ A diagnosis in $\mathcal{G}$. $Obs^{rec}$ Recall of observation.
$\mathcal{G}$ Procedural graphs. $Obs^{comp}$ Completeness of observation.
$g_{i}$ Procedural subgraph for disease $Exp^{com}$ Completeness of explanation.
$\mathcal{K}$ knowledge graphs. $Exp^{all}$ Completeness of all explanation.
$k_{i}$ knowledge subgraphs for disease $i$. $M$ An language model.
$\mathcal{P}_{i}$ Supporting edge collection for $k_{i}$. $\mathcal{E}$ Collection of annotated deductions.
$p$ A premise defined in $K$.
$\mathcal{F}_{i}$ Procedural edge collection for $g_{i}$.
$\mathcal{O}$ Collection of annotated observation.
$o$ An annotated observation.
$z$ Rationale for a deduction.
$d_{0}$ Root diagnosis for $g_{i}$.
Table 9: Notations defined in this paper.
A.2 Amended Data Points
Our proposed dataset aims to evaluate whether LLMs can provide a complete diagnostic reasoning process comparable to that of human doctors. To achieve this, we intended to select notes from the MIMIC database that contain comprehensive signs and symptoms as observations, enabling physicians to annotate the notes leading to a final PDD. For disease category like heart failure, MIMIC offers ample data, allowing us to choose notes with complete observations. However, for PDDs such as bacterial pneumonia, the number of relevant notes is limited, and many lack critical evidence necessary for diagnosis (e.g., sputum culture). We observed that in some notes, the section under the title ’sputum culture’ was left blank. We suspect that this might be due to some information being missed in MIMIC. To annotate such cases, we ask physicians add the necessary observations to support the diagnosis. In total, we made amendments to 73 notes. These notes all lacked evidence for a final PDD diagnosis, and in each note, only one observation was added as evidence. Thus, the modifications to the original content of the notes were minimal. For example, in a note where the PDD is bacterial pneumonia, we only added the following description under ’pertinent results’: ’Multiple organisms consistent with Haemophilus influenzae.’
To better illustrate the structure of our dataset and identify which data has been amended and what content has been added, we have provided a detailed CSV file on GitHub (https://github.com/wbw520/DiReCT/tree/master/utils/data_loading_analysisi). This file contains six columns, which record the following information: Disease Category, PDD, Data Root, Whether Amended, Amended Part, and Amended Content. The Data Root column records the path and filename of each note. We have stored the original note information and our annotations within a JSON file. The version submitted for review to PhysioNet follows the same storage format. In the Whether Amended column, notes that have been amended are marked as ’Yes,’ with the Amended Part and Amended Content columns specifying which part of the note was modified and what content was added. Additionally, we have provided several synthetic annotated samples (non-MIMIC data) on GitHub, along with detailed instructions on the format of the annotated data and how to parse each JSON file.
A.3 Structure of Knowledge Graph
@@
We first show the notations definition on Table 9. The entire knowledge graph, denoted as $\mathcal{K}$, is stored in separate JSON files, each corresponding to a specific disease category $i$ as $\mathcal{K}_{i}$. Each $\mathcal{K}_{i}$ comprises a procedural graph $\mathcal{G}_{i}$ and the corresponding premise $p$ for each disease. As illustrated in Figure 8, the procedural graph $\mathcal{G}_{i}$ is stored under the key "Diagnostic" in a dictionary structure. A key with an empty list as its value indicates a leaf diagnostic node as $d^{\star}$. The premise for each disease is saved under the key of "Knowledge" with the corresponding disease name as an index. For all the root nodes (e.g., Suspected Heart Failure), we further divide the premise into "Risk Factors", "Symptoms", and "Signs". Note that each premise is separated by ";".
Our knowledge graph was directly constructed by human physicians who followed authoritative diagnostic guidelines and incorporated their clinical experience. For Cardiology, Gastroenterology, Neurology, Pulmonology, and Endocrinology, the knowledge graph was built by 2, 1, 2, 2, and 1 specialists from the respective departments. The construction process involved first defining the procedural graph $g_{i}$ for each category, followed by supplementing $g_{i}$ with the detailed premises corresponding to each diagnosis d to build $k_{i}$. The complete knowledge graphs are available on GitHub (https://github.com/wbw520/DiReCT/tree/master/utils/data_annotation).
@@ B.1 Prompt Settings
In this section, we demonstrate the prompt we used for each module (From Table 10-12 for narrowing-down, perception, and reasoning module, respectively).
@@
In Table 12, {disease} is the disease category name and {disease_option} is consisted by the children nodes $\{d_{n}\}$. Similarly, the premise on the blue is only available for the $\mathcal{K}$ setting. It provides the premise that are criteria for the diagnosis of each children node. {observation} is the extracted $\hat{\mathcal{O}}$ in previous step. We provide all the prompts and the complete implementation code on GitHub.
B.2 Diagnostic Reasoning Under Conditions of Incomplete Observation
In real-world scenarios, doctors often have to make diagnoses based on incomplete information. To explore this, we conducted experiments on the 73 amended cases. One set of experiments used the unmodified original notes, labeled as "Original," while the other set used notes with added observations, labeled as "Amended." We tested three models—Llama3 70B, GPT-3.5-turbo, and GPT-4 turbo—under two settings: one with only the procedural graph $\mathcal{G}$ and the other with the complete knowledge graph $\mathcal{K}$. The results are presented in Tables 5 and 6. We can observe that in both $\mathcal{G}$ and $\mathcal{K}$ settings, the performance on the Amended data was consistently better across all metrics compared to the Original data. This suggests that even a single added observation can significantly impact the model’s diagnostic reasoning.
Additionally, we found that under the Amended data, using $\mathcal{K}$ led to both better diagnostic outcomes and improved explanability, aligning with the analysis in our paper. However, when using $\mathcal{K}$ on the Original data, while explanability improved, diagnostic accuracy actually decreased.
We conducted a detailed analysis of the 73 Original data’s results from GPT-4. We found that GPT-4 was still able to correctly deduce the final PDD in 24 cases using $\mathcal{G}$ and 19 cases using $\mathcal{K}$. This indicates that the model possesses some level of uncertain reasoning capability. However, upon further inspection, we found that in some cases, the model used completely irrational observations as evidence, such as directly using "cough" as evidence for diagnosing "bacterial pneumonia". Additionally, there were 7 cases using $\mathcal{G}$ and 13 cases using $\mathcal{K}$ where the reasoning stopped before the final PDD diagnosis. This suggests that the model recognized the lack of sufficient evidence to derive the PDD and adhered faithfully to the diagnostic knowledge graph. Moreover, using appeared to help the model better understand this limitation, however, decrease the accuracy. These results indicate that employing the knowledge graph acts more like a trade-off: using only $\mathcal{G}$ results in a higher tendency for uncertain reasoning, while using the full $\mathcal{K}$ makes the model more cautious.
Limitation of current implementation. Once a knowledge graph is provided, the focus shifts to whether the LLM follows the graph’s rules well. However, we consider the knowledge graph as an inferential framework rather than a set of rules. This framework provides decision-making paths for the LLM, but the LLM still needs to perform reasoning within it. Even when strictly following the knowledge graph, the LLM still needs to perform semantic analysis and context understanding in order to select the node that best suits the current situation among multiple possible paths (sub-nodes) in the knowledge graph. Therefore, the role of the LLM in this process is not merely to ’follow the rules,’ but to make logical path selections based on its understanding of the input data, which itself is a reflection of reasoning ability. This often requires the model or algorithm to consider previous steps in each stage of reasoning and to update observations accordingly. Even revise or backtrack the diagnosis step. However, our baseline method did not account for this and thus cannot fully exploit this capability of the LLM, which is a current limitation. We did try some designs to give those abilities to LLMs, such as providing previous steps of reasoning for the current stage as input prompts or update observations. However, even GPT-4 cannot show high instruction following ability to realize them (maybe the input is too long or prompt setting problems).
For evaluation, we used diagnostic processes annotated by human doctors as ground truth. Therefore, whether the KG is provided or not, the model’s output needs to align with the ground truth. Our dataset allows for evaluation with and without the KG, but our baseline method is not effective at handling scenarios without the KG (this is much more challenging). How to utilize and explore the LLM’s reasoning ability in this scenario is one of our future research directions.
B.3 Details of Automatic Evaluation
@@
For human evaluation, among the three specialists, two are from Cardiology and one is from Gastroenterology. Given that the notes originate from different medical domains, there is a possibility that the specialists may not be entirely accurate. However, this evaluation does not demand highly specialized knowledge, and it can be adequately covered by their expertise.
We also included an experimental result comparing the judgment differences between Llama3 8B and GPT-4 Turbo. The evaluation was performed on the diagnostic outcomes (across the entire dataset) from Llama3 70B and GPT-4 Turbo, using $\mathcal{G}$ as additional knowledge. We calculated the consistency rate for matching observations and the corresponding rationalization. As shown in Table 15, the differences in judgment between the two models are not obvious and are more consistent in observation discrimination. There are also some variations across different disease domains, with the highest similarity in observation discrimination found in Endocrinology, while the rationalization is most similar in Neurology.
Additionally, we provided results using GPT-4 Turbo for automatic evaluation, compared to those shown in Table 16 (which used Llama3 8B). The results indicate that GPT-4 Turbo tends to yield higher observation matching and more stringent rationalization discrimination. However, the largest difference does not exceed 5%. Considering the cost of GPT-4, Llama3 8B is a more efficient option.
Table 15: Judgement consistency between LLama3 8B and GPT-4 turbo.
LLama3 70B GPT-4 turbo
Domain Observation Rationalization Observation Rationalization
Cardiology 0.885±${}_{\pm\text{0.095}}$ 0.761±${}_{\pm\text{0.268}}$ 0.827±${}_{\pm\text{0.146}}$ 0.861±${}_{\pm\text{0.273}}$
Gastroenterology 0.862±${}_{\pm\text{0.088}}$ 0.676±${}_{\pm\text{0.361}}$ 0.810±${}_{\pm\text{0.167}}$ 0.755±${}_{\pm\text{0.316}}$
Neurology 0.846±${}_{\pm\text{0.090}}$ 0.831±${}_{\pm\text{0.211}}$ 0.856±${}_{\pm\text{0.124}}$ 0.963±${}_{\pm\text{0.106}}$
Pulmonology 0.808±${}_{\pm\text{0.131}}$ 0.703±${}_{\pm\text{0.317}}$ 0.786±${}_{\pm\text{0.152}}$ 0.779±${}_{\pm\text{0.287}}$
Endocrinology 0.911±${}_{\pm\text{0.104}}$ 0.783±${}_{\pm\text{0.304}}$ 0.868±${}_{\pm\text{0.145}}$ 0.793±${}_{\pm\text{0.340}}$
Overall 0.869±${}_{\pm\text{0.102}}$ 0.734±${}_{\pm\text{0.321}}$ 0.838±${}_{\pm\text{0.144}}$ 0.806±${}_{\pm\text{0.305}}$
Table 16: Result of using GPT-4 turbo and LLama3 8B for automatic evaluation.
Observation Explanation
Judgement Models $\textit{Obs}^{\text{pre}}$ $\textit{Obs}^{\text{rec}}$ $\textit{Obs}^{\text{comp}}$ $\textit{Exp}^{\text{com}}$ $\textit{Exp}^{\text{all}}$
GPT-4 turbo LLama3 70B 0.317±${}_{\pm\text{0.161}}$ 0.576±${}_{\pm\text{0.195}}$ 0.294±${}_{\pm\text{0.159}}$ 0.348±${}_{\pm\text{0.300}}$ 0.107±${}_{\pm\text{0.118}}$
GPT-4 turbo 0.465±${}_{\pm\text{0.190}}$ 0.514±${}_{\pm\text{0.157}}$ 0.408±${}_{\pm\text{0.201}}$ 0.437±${}_{\pm\text{0.335}}$ 0.187±${}_{\pm\text{0.191}}$
LLama3 8B LLama3 70B 0.277±${}_{\pm\text{0.146}}$ 0.537±${}_{\pm\text{0.192}}$ 0.256±${}_{\pm\text{0.142}}$ 0.395±${}_{\pm\text{0.320}}$ 0.112±${}_{\pm\text{0.110}}$
GPT-4 turbo 0.446±${}_{\pm\text{0.207}}$ 0.491±${}_{\pm\text{0.180}}$ 0.371±${}_{\pm\text{0.186}}$ 0.475±${}_{\pm\text{0.363}}$ 0.199±${}_{\pm\text{0.181}}$
@@
B.4 More Prediction Samples
10 shows another sample generated by GPT-4. The ground-truth PDD of the input clinical note is and Heart Failure (HF). In these figures, purple, orange, and red indicate explanations only in the ground truth, only in prediction, and common in both, respectively; therefore, red is a successful prediction of an explanation, while purple and orange are a false negative and false positive.
In Figure 10, the model’s predictions do not align well with those of a human doctor. Key observations, such as the relationships between BNP and LVEF, are incorrectly identified, leading to a final misdiagnosis.
B.5 Experiments for No Extra Knowledge
We demonstrate the prompt used for $\mathcal{D}^{\star}$ and no knowledge settings in Table 17 and Table 18, respectively. {note} is the text of whole clinical note and {disease_options} in Table 17 is the name of all leaf node $\mathcal{D}^{\star}$.
B.6 Experimental Settings